OpenStudy (anonymous):

r=2/3t +v...solve for t...thanks

OpenStudy (anonymous):

r=2/3t + v. Multiply both sides by 3. 3r=2t+3v Now divide both sides by 2. 3/2r = t + 3/2v 3/2r-3/2v=t Hope this helps. Thanks, Sean Walsh Facebook: Tutor Sean http://www.tutorsean.net

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

NP. Anytime. Love helping others. That is why I made a website with free math videos and a facebook community for math and other topics. :)